Circular No. 8848

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V2467 CYGNI S. Mazuk, D. K. Lynch, R. J. Rudy, R. W. Russell, R. L. Pearson,

The Aerospace Corporation; C. E. Woodward, University of Minnesota; R. C. Puetter, Center for Astrophysics and Space Science, University of California at San Diego; and R. B. Perry, Langley Research Center, NASA, report 0.4- to 2.5-μm spectroscopy of this nova (cf. IAUC 8821) using the Lick 3-m telescope (+ VNIRIS) on May 7 UT, as well as 0.8- to 5.5-μm spectroscopy using the Infrared Telescope Facility 3-m telescope (+ SpeX) on May 31. The spectrum increased in excitation between the two epochs, displaying coronal lines of [Si VI], [Si VII], [Ca VIII], [S VIII], and [S IX] at the later date. All coronal lines had complex profiles. He II emission was also present, in addition to the unidentified novae lines (Rudy et al. 2002, Ap.J. 573, 794) at 1.11, 1.19, 1.55, and 2.10 μm. Despite the presence of these high-excitation features, emission from C I, N I, and O I persisted. There was no evidence of dust formation. The interstellar reddening derived from the O I lines was E(B − V) = 1.5.

COMETS C/2007 J1–J6 (SOHO) Additional near-sun comets (cf. IAUC 8847) have been found on SOHO

website images — all being Kreutz sungrazers except for C/2007 J1 (Meyer group). C/2007 J1 was stellar in appearance and reached mag ∼ 6. C/2007 J5 was fuzzy, very faint, and tailless. C/2007 J6 was very diffuse and somewhat elongated, with a hint of a tail. The remaining three objects were very small, slightly diffuse, and faint (mag ∼ 7.5–8).

Comet 2007 UT α_{2000} δ_{2000} Inst. F MPEC
C/2007 J1 May 2.579 2^{h}43.^{m}5 +15^{o}29′ C3/2 RK 2007-L03
C/2007 J2 3.121 2 44.8 +14 17 C2 BZ 2007-L03
C/2007 J3  4.746 2 54.4 +14 17 C3/2 TH 2007-L03
C/2007 J4  4.771 2 50.9 +14 43 C2  TH 2007-L03
C/2007 J5  7.538 3 01.3 +15 25 C2  BZ 2007-L16
C/2007 J6  9.146 3 06.3 +15 48 C2  BZ 2007-L16

COMET 8P/TUTTLE C. W. Hergenrother reports his recovery of this comet as a stellar

object of mag 20.3–20.4 in co-added CCD exposures taken on Apr. 22.5 and 26.5 UT with the Catalina 1.54-m reflector. The astrometry, revised orbital elements, and an ephemeris appear on MPEC 2007-L45.

2007 June 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green